OpenStudy (nicole143):

@mathmale Hi again, I'm working out how to find f/g and it's domain. I think I got f/g but I'm confused on if it's domain is "x is NOT equal to 4" or "x is Not equal to -4". Can you explain how to get this?

OpenStudy (mathmale):

Hi, Nicole! Happy to "see" you again. If the domain is already given as "x is not equal to 4" or "x is not equal to -4," I could say with certainty that the composite function f/g has a denominator that is zero for x=4 or x=-4. Let's discuss this until you feel comfortable with the concepts involved.

OpenStudy (nicole143):

Okay, this is what I have done.

OpenStudy (mathmale):

This image helps so much!!

OpenStudy (nicole143):

Good! (:

OpenStudy (mathmale):

We can get the domain of f/g directly from the algebraic expression BEFORE factoring and simplification. \[\frac{ x ^{2}-16 }{ x+4 }\] (ob viously) has the denominator x+4. Let this x+4=0 and solve for x. THAT is the value that x can NOT have.

OpenStudy (mathmale):

Your factoring and reduction, while done correctly, doesn't help much with finding the domain.

OpenStudy (nicole143):

Okay, so how do I find if it will either be a -4 or +4?

OpenStudy (mathmale):

I suspect you have a rational function whose denominator is (x+4)(x-4). Am I right? If so, set that denominator, (x+4)(x-4), = to zero and solve for x. One solution is x=4; the other is x=-4. Thus, the domain of that function is

OpenStudy (mathmale):

"all real numbers other than 4 and -4"

OpenStudy (mathmale):

or, alternatively, (-infinity,-4) U (-4, 4) U (4, +infinity).

OpenStudy (nicole143):

I can only have one though.. It's either all but -4 or all but +4. I may not be understand though, I may be thinking about this differently and the answer is right there..

OpenStudy (mathmale):

Back to the drawing board. If those are the two choices, then the denominator of your f/g is probably (x+4). If it is (x+4), then the domain of your f/g is "all real numbers other than -4." If it's (x-4), then the domain of your f/g is "all real numbers other than 4."

OpenStudy (nicole143):

Okay, so if I come down to the answer x = 4 then its domain is going to be all real numbers except -4?

OpenStudy (mathmale):

Not quite. let me re-phrase your question: "OK, so if I have (x+4) in the denominator of my f/g, then the domain of f/g is going to be "all real numbers other than -4."

OpenStudy (mathmale):

Note: you said, "x=4" I said "(x+4)".

OpenStudy (nicole143):

Oh, I didn't know they had to have the ( ) -my answers choices don't have them so I figured that they weren't needed. May I ask another simple quick yes or no question in this same post?

OpenStudy (mathmale):

Sure, please go right ahead!

OpenStudy (nicole143):

When factoring out an equation -Example: (5x + 2)(2x + 1)- would you have 10x + 5x + 4x + 2 or 10x^2 + 5x + 4x^2 + 2??

OpenStudy (mathmale):

Nicole, I'm quite certain you mean, "When multiplying out a function expressed as the product of factors, such as (5x+2)(2x+1), which of the following would be the result? ... " Answer: 10x^2+9x+2. Mind trying that again? Hint: x*x = x^2.

OpenStudy (nicole143):

That's all I needed to know - I had x and not x^2 - And I thought "That doesn't look right" and I meant using "foil" not factoring.

OpenStudy (mathmale):

So glad you feel good about this problem and hopefully about the last one as well. Again, I have to get off my computer (after having spent much of the day sitting in front of it), so I'll say good night and welcome you to come back again soon! Great working with you! MM

OpenStudy (nicole143):

Goodnight! Thank you!

OpenStudy (mathmale):

Bye, Nicole!